OpenStudy (chrisplusian):

Can someone help me with a double integral? Its been a long time and I forgot

OpenStudy (chrisplusian):

\[\int\limits_{0}^{\pi/2}dy \int\limits_{0}^{5}xdx\]

OpenStudy (anonymous):

What's the problem?

OpenStudy (chrisplusian):

\[\int\limits_{0}^{\pi/2}(dy) \int\limits_{0}^{5}(xdx)\]

OpenStudy (chrisplusian):

did it show up that time?

OpenStudy (anonymous):

Yes, it's there, but where exactly do you need help?

OpenStudy (anonymous):

Just integrate respect to dx and dy separately, remember integrating a constant will give you a variable, and integrating just a variable is (x^(n+1))/(n+1)

OpenStudy (anonymous):

\[\large y |_{0 \rightarrow \pi/2} \times \frac{ x^2 }{ 2 } |_{0 \rightarrow 5}\]

OpenStudy (chrisplusian):

I apologize the site was giving me a problem for a few minutes

OpenStudy (anonymous):

Oh no worries, it's spherical coordinates

OpenStudy (chrisplusian):

I did just what your saying but I did not get the same result as my textbook

OpenStudy (anonymous):

Well, I have to go but I can tag someone that can help you @ganeshie8

OpenStudy (chrisplusian):

yes but everything is laid out, I understand how to set it up but I am not getting the same result as them

OpenStudy (chrisplusian):

could you just check to see if you get the same result as the book?

OpenStudy (chrisplusian):

I got \[\frac{ 25\pi }{ 4 } = 19.63\]

OpenStudy (chrisplusian):

@ganeshie8

OpenStudy (chrisplusian):

can you help please?